OpenStudy (anonymous):

Evaluate0.5g-g2 when g=4 -14 4 18 -6

OpenStudy (anonymous):

It can be written as: 1/2g - g^2 So when g = 4, 1/2 * 4 - (4^2) = 2 - 16 = - 14

OpenStudy (anonymous):

Whats the answer